# QUIDS: Query Intent Generation via Dual Space Modeling

### v2

@@ 6. Conclusion
In this paper, we introduced a novel dual-space modeling approach for the task of query intent generation. Our approach implements contrastive learning in both the encoding and decoding phases, combined with intent-driven hard negative augmentation during data preprocessing, to automatically generate detailed and precise intent descriptions. Experimental results show that our model effectively filters out irrelevant information from the relevant intent space, leading to more accurate intent descriptions and greater transparency in the retrieval process. For future work, we aim to extend our approach to conversational search by first mining exploratory needs and then explaining the understanding of query intents.
Acknowledgements.
@@
Appendix A Appendix A
A.1. Implementation details for baseline models
@@
A.2. Correlation with Human Evaluation
Table 5. Spearman ($\rho$) and Kendall-Tau ($\tau$) correlations between Human evaluation and LLM evaluation of different metrics.
Correlation Model Fluen. Align. Inclu. Exluc.
$\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$
Corr (Human, GPT-4o) SegEnc 0.377 0.304 0.409 0.310 0.184 0.159 0.369 0.306
QUIDS 0.476 0.375 0.517 0.412 0.252 0.199 0.361 0.286
Corr (Human, LLaMa3.1) SegEnc 0.243 0.187 0.459 0.344 0.143 0.111 0.224 0.177
QUIDS 0.343 0.278 0.553 0.423 0.325 0.248 0.158 0.126
A.3. Breakdown Analysis on Intent Types
Table 6. Comparision of Human and LLM evaluation on informational and exploratory intent types on our model.
Method Intent Fluen. Align. Inclu. Exluc.
Human Info. 4.71 3.51 3.98 4.73
Expl. 4.85 4.04 3.88 4.74
@@ A.3. Breakdown Analysis on Intent Types
Figure 5. Boxplot of evaluation scores on 4 metrics of our model on different intent types.
[…]
@@
In Table 3, we observe a human preference over the SegEnc model on metric Factual Alignment, which measures how well the generated query intent description is factually aligned with the ground truth intent. We guess it is due to the model performance difference on different sub-datasets, or on different intent types. And hence we further analyse the human evaluations on different intent types. Table 6 present human and LLM evaluations on our model regarding two intent types. While humans prefer our model for exploratory intent with 4.04 (QUIDS) vs. 3.81 (SegEnc), SegEnc is favored for informative intent with 3.51 (QUIDS) vs. 3.68 (SegEnc). Since informative intent queries dominate, this leads to a lower average score for our model. We also present the evaluation scores distribution on different intent types over 50 the same test samples in subsection 5.4. Figure 5 is a breakdown of Figure 3 on our model for four metrics. Based on Table 6 and Figure 5, we have following insights:
  • Exploratory intent types generally outperform informational intent types across all metrics, except for human evaluations on inclusion. This finding, based on 50 test samples from both human and LLM evaluations, is consistent with the automatic evaluation results for the full test dataset, as shown in Table 4.
  • The boxplots for informational intent on the right side of each metric in Figure 5 exhibit a distribution similar to the overall performance of our model presented in Figure 3, suggesting that informational intent queries dominate the dataset and largely drive performance outcomes. However, exploratory intent queries outperform in this particular task, indicating their stronger performance despite being less frequent.
A.4. Breakdown Analysis by document length
Figure 6 shows the results of the automatic evaluation, broken down based on document length.
Figure 6. Comparison of automatic evaluation on our model QUIDS on different input lengths.
[…]
Appendix B Appendix B
B.1. Evaluation prompt on Fluency
@@
Assign a rating from 1 to 5 based on the level of fluency.
@@
B.2. Evaluation prompt on Factual Alignment
Your task is to rate the query intent description on one metric.
@@ B.2. Evaluation prompt on Factual Alignment
Query:
{{Query}}
Relevant documents:
{{Relevant documents}}
Irrelevant documents:
{{Irrelevant documents}}
Generated Intent:
{{Generated intent}}
Ground Truth Intent:
{{Gound truth intent}}
@@
B.3. Evaluation prompt on Inclusion Score
Your task is to rate the query intent description on one metric.
@@ B.3. Evaluation prompt on Inclusion Score
Query:
{{Query}}
Relevant documents:
{{Relevant documents}}
Irrelevant documents:
{{Irrelevant documents}}
Generated Intent:
{{Generated intent}}
Ground Truth Intent:
{{Gound truth intent}}
@@
B.4. Evaluation prompt on Exclusion Score
Your task is to rate the query intent description on one metric.
@@ B.4. Evaluation prompt on Exclusion Score
Query:
{{Query}}
Relevant documents:
{{Relevant documents}}
Irrelevant documents:
{{Irrelevant documents}}
Generated Intent:
{{Generated intent}}
Ground Truth Intent:
{{Gound truth intent}}

### v3

@@ 6 Conclusions
We introduced a novel dual-space modeling approach for the query intent generation task. Our approach implements contrastive learning in both encoding and decoding phases, combined with intent-driven hard negative augmentation during data preprocessing, to automatically generate detailed and precise intent descriptions, surfacing what the system likely inferred the query to mean. Experimental results show that our model can effectively filter out irrelevant information from the relevant intent space, leading to more accurate intent descriptions than all baselines, including models for Query-Focused Summarization. In future work, we plan to improve contextual understanding in distinguishing relevant from irrelevant information and extend our approach to conversational search by mining exploratory needs and explaining the understanding of query intents. Our long-term aim is to improve the transparency in the retrieval process, in particular for exploratory search needs.
Acknowledgments
@@
Limitations
Training Efficiency Trade-offs. Augmenting irrelevant documents enhances robustness but linearly increases training time. We therefore limited negative documents to three per query, which partially alleviates this trade-off but remains suboptimal. To fundamentally resolve this efficiency bottleneck, two promising directions are: (1) adaptive dynamic sampling that prioritizes high-impact negatives through real-time gradient analysis, and (2) curriculum-based augmentation progressively introducing harder negatives as training stabilizes.
Dataset Imbalance. Informational queries dominate the training data over exploratory ones. While our model shows promising performance in exploratory search scenarios, this bias limits deeper intent analysis. Future work should expand out experiments to more datasets, focusing on exploratory queries. One option would be to use LLM-generated synthetic data, specifically creating pseudo-documents that mimic multi-faceted exploratory intents. This approach maintains intent modeling consistency while enabling systematic investigation of query complexity, without requiring manual annotation efforts.
Appendix A Detailed Comparison with Related Work
A detailed comparison between our task—Query intent description generation —and related tasks such as Query Understanding (QU), Query-Focused Summarization (QFS), and Pseudo-Relevance Feedback (PRF) is provided in Table 6.
Appendix B Experimental Setting Details
B.1 Dataset Details
In constructing the Q2ID dataset, documents with multi-graded relevance labels were converted into binary labels, indicating whether each document is relevant to the query. The dataset is composed of:
  • 510 entries from TREC tracks (Dynamic Domain 2015–2017, Robust 2004)
  • 4,878 entries from SemEval-2015/2016 Task 3 on Community Question Answering
Each data point is formatted as a quadruple: <query, relevant documents, irrelevant documents, intent description>.
The average query length is 7.2 tokens, and the average intent description length is 45.5 tokens. We follow the original split of Q2ID: 5,000 queries for training, 100 for validation, and 258 for testing.
B.2 Sensitivity Analysis of Loss Weights
We conduct a sensitivity analysis through a grid search over different combinations of the loss weights ($\lambda_{0}$, $\lambda_{1}$, $\lambda_{2}$). The results of several representative configurations are summarized in Table 5.
$\lambda_{0}$ $\lambda_{1}$ $\lambda_{2}$ RG-1 RG-2 RG-3 BS
- - - 34.53 14.42 26.85 63.33
0.33 0.33 0.33 33.36 14.64 26.22 63.45
0.20 0.20 0.60 34.51 15.36 27.13 63.52
0.30 0.30 0.40 34.17 14.31 26.30 63.41
0.10 0.10 0.80 32.37 14.59 25.81 62.87
0.20 0.40 0.40 33.20 14.56 26.10 63.22
Table 5: Sensitivity analysis of loss weights. ‘-’ indicates that the $\lambda$ parameters were treated as learnable during training.
We observed that configurations assigning relatively higher weight to $\lambda_{2}$ tend to yield better overall performance. Overly large $\lambda_{2}$ values (e.g., 0.8) degrade performance by reducing the contribution of the other loss components. Configurations that assign more weight to $\lambda_{0}$ or $\lambda_{1}$ alone lead to inferior performance, sometimes resulting in no valid generation, and are therefore omitted for clarity. Based on these observations, we adopt $(\lambda_{0},\lambda_{1},\lambda_{2})=(0.2,0.2,0.6)$ as the default setting.
Task Query Understanding (Classification / Clustering / Expansion) Query-Focused Summarization (QFS) Pseudo-Relevance Feedback (PRF) Our Task: Query Intent Generation (QIG)
Goal Predict query intent classes, discover latent topics, or expand queries for better retrieval performance. Summarize relevant documents to help users consume content. Refine or reformulate queries to improve retrieval performance. Generate a natural language description of the search system’s inferred intent behind a query.
Output Form Labels (e.g., informational/navigational), clusters, or expanded query terms. Natural language summary (abstractive or extractive). Modified query or re-weighted terms. Natural language explanation of inferred query intent.
Use of Irrelevant Documents Not used. Focus is on query-only or top-ranked documents. Rarely used; mainly uses pseudo-relevant documents. Not used; PRF assumes top-ranked documents are relevant. Explicitly contrasts relevant and irrelevant documents for intent disentanglement.
Application Stage Pre-retrieval; typically before document scoring. Post-retrieval summarization. Interleaved or pre-retrieval (used for re-ranking or expansion). Post-retrieval; supporting user query refinement and retrieval debugging
User Utility Improves ranking accuracy and personalization; not visible to users. Helps users consume content more efficiently. Improves recall or relevance through backend query rewriting. Helps users understand potential mismatches between their intended query meaning and the system’s inferred intent.
Table 6: Comparison between Query Understanding (QU) tasks, Query-Focused Summarization (QFS), Pseudo-Relevance Feedback (PRF), and our Query Intent Generation (QIG) task.
B.3 Implementation
To balance efficiency and effectiveness during IDNA augmentation, we set the expected number of irrelevant documents per query to three for training. This results in augmenting 1,984 queries with at least three irrelevant documents each. Training is conducted for 10 epochs using the Adam optimizer with a learning rate of 0.0001. During decoding, we set a maximum sequence length of 256 tokens and apply beam search with a beam size of 4. We also set a no-repeat n-gram size of 3 to reduce redundancy.
B.4 Baselines
Large Language Model (LLM) baseline: LLaMa3.1-8B-Instruct (AI, 2024): We evaluate the LLaMa3.1-8B instruction-tuned text-only model under both zero-shot and two-shot settings and conduct five experimental runs for each setting. For two-shot setting, we randomly using two different examples per run – one sourced from TREC and the other from SemEval. OpenAI o3 (OpenAI, 2025): We evaluate the reasoning-focused OpenAI o3 model. While the model internally generates reasoning tokens, we focus our analysis on the visible output tokens. The same experimental settings and number of runs are used for comparison.
Query Focused Summarization (QFS) baselines: RelReg (Vig et al., 2022) and RelRegTT (Vig et al., 2022): two-step approaches for QFS consisting of an score-and-rank extractor and an abstractor. The extractor is trained to predict ROUGE relevance scores and then the ranked results based on ROUGE are passed to the abstractor. SegEnc (Vig et al., 2022): an end-to-end approach tailored for handling longer input texts. SegEnc splits a long input into fixed-length overlapping segments and encodes them separately. The encoding sequences are concatenated so that the decoder can attend to all encoded segments jointly. Socratic (Pagnoni et al., 2023): an unsupervised, question-driven pretraining approach designed to tailor generic language models for controllable summarization tasks. Qontsum (Sotudeh and Goharian, 2023): an abstractive summarizer that applied Generative Information Retrieval (GIR) techniques. It builds on SegEnc by adding a segment scorer and contrastive learning modules. We train the QFS baselines on the Q2ID dataset using the original code provided by the authors, except for Qontsum, which we independently reproduced.
B.5 Implementation Details for Baselines
@@
Correlation Model Fluency Alignment Inclusion Exclusion Average
$\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$
Corr (Human, GPT-4o) SegEnc 0.320 0.252 0.387 0.295 0.158 0.121 0.375 0.309 0.310 0.244
QUIDS 0.476 0.375 0.495 0.386 0.265 0.199 0.361 0.286 0.399 0.312
Corr (Human, LLaMa3.1) SegEnc 0.224 0.173 0.434 0.329 0.153 0.116 0.343 0.272 0.289 0.223
QUIDS 0.373 0.294 0.557 0.424 0.365 0.262 0.158 0.126 0.363 0.276
Table 7: Spearman ($\rho$) and Kendall-Tau ($\tau$) correlations between Human evaluation and LLM evaluation of different metrics.
Method Intent Fluen. Align. Inclu. Exclu.
Human Info. 4.78 3.71 4.04 4.80
Expl. 4.89 4.19 4.14 4.81
@@
Table 8: Comparison of Human and LLM evaluation on informational and exploratory intent types on our model.
Appendix C Correlation with Human Evaluation
Appendix D Evaluation on Intent Types
@@
In Table 3, we observe a human preference over the SegEnc model on metric Factual Alignment, which measures how well the generated query intent description is factually aligned with the ground truth intent. We guess it is due to the model performance difference on different sub-datasets, or on different intent types. And hence we further analyse the evaluations on different intent types.
Automatic Evaluation
In the 258 test samples, there are 20 queries with exploratory intents and 238 with informational intents. As shown in Table 4, queries with exploratory intents substantially outperform those with informational intents, achieving 60% higher ROUGE-2 scores and 20% higher BERTScores. This indicates that our model is better suited for exploratory queries. This finding contrasts with the results of (Zhang et al., 2020), where the CtrsGen model performed slightly better on the informational SemEval queries than on the exploratory TREC queries. A potential explanation is that the backbone language model used in our approach more freely generates text than the GRU model used in (Zhang et al., 2020), particularly when reconstructing complex scenarios for informational intents. This is an aspect that makes our approach more suitable for exploratory search rather than informative search.
Human and LLM Evaluation
Table 8 presents human and LLM evaluations on our model regarding two intent types. In general, exploratory intents consistently outperform informational intents across all metrics. This finding, derived from 50 test samples, aligns with automatic evaluation results on the full test dataset (Table 4). Figure 4 shows the evaluation score distribution by intent type on our model, compared to the overall model performance in Figure 3. The informational intent distribution closely mirrors the overall performance, suggesting that informational queries dominate the dataset and largely influence performance. However, exploratory queries, despite being less frequent, demonstrate superior performance in this task. When diving into the factual alignment in Figure 3, while humans prefer our model for exploratory intent with 4.19 (QUIDS) vs. 3.94 (SegEnc), SegEnc is favored for informative intent with 3.71 (QUIDS) vs. 3.89 (SegEnc). Since informative intent queries dominate, this leads to a lower average score for our model on this metric. These findings indicate that our model is well-suited for exploratory search.
Appendix E Failure Case Study
Figure 7: Failure case study. Token-level decoder cross-attention weights for a generated intent token (red) are shown with (c) and without (d) an irrelevant document. Deeper color indicates a higher value.
Figure 7 illustrates a failure example of filtering irrelevant information when an irrelevant document is provided. The token-level decoder cross-attention weights are compared when generating a content word in the intent, with (c) and without (d) an irrelevant document. When generating the keyword ‘UK’ and ‘Dubai’, the model mainly focuses on ‘petrol’ and ‘cigarettes’ in the relevant documents for both (c) and (d), which are also contextually important in the generated intent. However, the model fails to recognize the relationship between ‘middle east’ in the irrelevance document and ‘Dubai’, leading to the unwanted inclusion of ‘Dubai’ in the intent 2. This highlights that our model may struggle with excluding information that requires commonsense reasoning or domain-specific knowledge. A direction for future work is to develop advanced approaches that enhance contextual understanding for complex scenarios.
Appendix F LLM-based Evaluation Details
Following the method of (Liu et al., 2023b), we use LLaMa3.1-8B-Instruct [^3] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct. and GPT-4o [^4] https://openai.com/index/hello-gpt-4o/. as instruction-tuned evaluators to assess the generated intent across four qualitative metrics. Specifically, we define the evaluation task and criteria, prompting the LLM to generate chain-of-thoughts (CoT) for the ‘Evaluation Steps’. For LLaMa3.1-8B-Instruct, we use the output token probabilities from the LLMs to normalize the scores and take their weighted summation as the final results:
$score=\sum_{i=1}^{n}p(s_{i})\times s_{i}$ (11)
where $S=\{s_{1},s_{2},...,s_{n}\}$ represents the predefined score set from the prompt, with a maximum value of 5 in our case. For the close-sourced GPT-4o, we sample 20 times to estimate the token probabilities. An example prompt for each model is presented below.
F.1 General Evaluation Prompt
@@
<<EvaluationSteps>>
@@
- <<MetricName>>:
F.2 Evaluation Prompt on Fluency
Evaluation Criteria:
Evaluation steps:
Assign a rating from 1 to 5 based on the level of fluency.
Evaluation Form (scores ONLY):
@@ F.2 Evaluation Prompt on Fluency
F.3 Evaluation Prompt on Factual Alignment
@@
F.4 Evaluation Prompt on Inclusion Score
@@
F.5 Evaluation Prompt on Exclusion Score